OpenStudy (anonymous):

Write the point-slope equation for the line with the given slope that contains the given point. m=-4/-3 ; (7, -1)

OpenStudy (anonymous):

the slope form for this is Y= m(y-1) (x-1) or you could use MX+B form as well. does this help?

hartnn (hartnn):

lets use the form y=mx+c slope m= 4/3 given now u need to find c only you have a point (7,-1) on it so put x=7 and y=-1 in the equation and find the value of c got this? if yes,can u tell me what value u got of c?

OpenStudy (anonymous):

You just made my head spin.

OpenStudy (anonymous):

hey hartnn, Point slope form is what I typed right? I thought it he was asking to write the the point slope form? did I miss something here??

hartnn (hartnn):

ohh! slope point form,i missed that the slope point form is y-y1=m(x-x1) you have the values of m as 4/3 and (x1,y1)=(7,-1) just put the values,and u get the equation

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

y+1=-4/-3(x-7)